OpenStudy (anonymous):

The point P is on the unit circle. Find P(x, y)The y-coordinate of P is -1/5 and the x-coordinate is positive. from the given information.

OpenStudy (whpalmer4):

Any point on the unit circle satisfies the constraint \[x^2+y^2 = r^2\] where \(r = 1\) or\[x^2+y^2=1\] If you know \(y = -\dfrac{1}{5}\) then \[x^2 + (-\frac{1}{5})^2 = 1\]Can you solve that for \(x\)? You'll take the positive solution, as you've been told the x-coordinate is positive.